OpenStudy (anonymous):

Solve the problem by using a system of equations? The sum of two numbers is 53, and their difference is 21. Find the numbers.

jimthompson5910 (jim_thompson5910):

The sum of two numbers is 53 -----> x+y = 53

jimthompson5910 (jim_thompson5910):

and their difference is 21 ----> x - y = 21 this is making 'x' the larger number

jimthompson5910 (jim_thompson5910):

Add the two equations x+y = 53 x-y = 21 ---------- 2x+0y = 74 2x = 74 x = ??? Now use this solution of x to find the solution for y.

OpenStudy (anonymous):

x = 37 and y = 16

OpenStudy (anonymous):

thanks! I didn't know you had to add the 2 equations together

jimthompson5910 (jim_thompson5910):

that's just one of the many ways to do it

OpenStudy (anonymous):

as long as its the easiest/most efficient way. It seems pretty good to me